OpenStudy (anonymous):

How many solutions does the equation have? 4(2x+3)=2(3x-4) A. one solution B. infinite solutions C. no solution please hellpp @gorv @ganeshie8

OpenStudy (gorv):

divide by 2 on both side

OpenStudy (gorv):

2 on right side will get cancelled

OpenStudy (gorv):

2(2x+3)=3x-4

OpenStudy (gorv):

now take 2 inside the bracket

OpenStudy (amistre64):

you have 2 lines with different slopes as well ....

OpenStudy (gorv):

4x+6=3x-4

OpenStudy (gorv):

now subtract 3x from both side

OpenStudy (gorv):

and tell me what you got

OpenStudy (anonymous):

i have no idea how to do that

OpenStudy (anonymous):

so its one solution?

OpenStudy (texaschic101):

do you know how to distribute through the parenthesis ?

OpenStudy (gorv):

4x+6-3x=3x-4-3x

OpenStudy (gorv):

dont run for ans plzzz ..u r not goin to learn anything thn

OpenStudy (gorv):

now tell me what u got after this equation??

OpenStudy (anonymous):

this is so hard idk umm7x

OpenStudy (anonymous):

@gorv

OpenStudy (gorv):

this similar like numbers just keep x aside and subtract term with x like number 4x-3x consider them without x and after subtracting just attach x with the result

OpenStudy (gorv):

so 4x-3x=????

OpenStudy (anonymous):

is it one solution just tell me i think its 1 solution

OpenStudy (gorv):

ohh u need just ans

OpenStudy (anonymous):

yes

OpenStudy (gorv):

lolll we dont give direct ans here